Please show formulas used. You're planning to save for his retirement 34 years from now. You plan to invest $4,200 per year for the first 7 years, $6,900 per year for the next 11 years, and $14,500 per year for the following 16 years. You make all investments at the end of each year. The investments will earn an annual rate of return of 9.7%. What will your investment be worth 34 years from now?
Answer #1
Step-1:Value of investment of first 7 years at the end of 34 years from now
Future value = Annual cash flow * Future value of annuity of 1 for 7 years * Future value of 1 for 27 years
= $          4,200.00 * 9.400178614 * 12.17808342
= $    4,80,799.87
Working:
Future value of annuity of 1 = (((1+i)^n)-1)/i Where,
= (((1+0.097)^7)-1)/0.097 i = 9.70%
= 9.400178614 n = 7
Future value of 1 for 27 years = (1+i)^n Where,
= (1+0.097)^27 i = 9.70%
= 12.17808342 n = 27
Step-2:Value of investment of next 11 years at the end of 34 years from now
Future value = Annual cash flow * Future value of annuity of 1 for 11 years * Future value of 1 for 16 years
= $          6,900.00 * 18.23382362 * 4.398514634
= $    5,53,392.01
Working:
Future value of annuity of 1 = (((1+i)^n)-1)/i Where,
= (((1+0.097)^11)-1)/0.097 i = 9.70%
= 18.23382362 n = 11
Future value of 1 for 27 years = (1+i)^n Where,
= (1+0.097)^16 i = 9.70%
= 4.398514634 n = 16
Step-3:Value of investment of next 16 years at the end of 34 years from now
Future value = Annual cash flow * Future value of annuity of 1 for 11 years
= $       14,500.00 * 35.03623334
= $    5,08,025.38
Working:
Future value of annuity of 1 = (((1+i)^n)-1)/i Where,
= (((1+0.097)^16)-1)/0.097 i = 9.70%
= 35.03623334 n = 16
Step-4:Value of investment 34 years from now
Investment worth 34 years from now = $    4,80,799.87 + $ 5,53,392.01 + $ 5,08,025.38
= $ 15,42,217.26